### v1

# Accelerating Residual Reinforcement Learning with Uncertainty Estimation

Abstract
Residual Reinforcement Learning (RL) is a popular approach for adapting pretrained policies by learning a lightweight residual policy that provides corrective actions. While Residual RL is more sample-efficient than finetuning the entire base policy, existing methods struggle with sparse rewards and are designed for deterministic base policies. We propose two improvements to Residual RL that further enhance its sample efficiency and make it suitable for stochastic base policies. First, we leverage uncertainty estimates of the base policy to focus exploration on regions in which the base policy is not confident. Second, we propose a simple modification to off-policy residual learning that allows it to observe base actions and better handle stochastic base policies. We evaluate our method with both Gaussian-based and Diffusion-based stochastic base policies on tasks from Robosuite and D4RL, and compare against state-of-the-art finetuning methods, demo-augmented RL methods, and other residual RL methods. Our algorithm significantly outperforms existing baselines in a variety of simulation benchmark environments. We also deploy our learned polices in the real world to demonstrate their robustness with zero-shot sim-to-real transfer.
Figure 1: We propose two improvements to accelerate Residual RL : 1) We use an uncertainty metric to constrain exploration by only adding a residual for uncertain base actions. 2) We modify the off-policy critic to learn the $Q$ function for the combined action.
I Introduction
Residual Reinforcement Learning is popular in robotics to improve the performance of pretrained policies by training a separate policy that outputs residual actions  [1, 2]. Directly fine-tuning the pretrained policy is often computationally expensive, especially in the case of policies with a large number of parameters [3, 4] and is prone to instability [5]. In contrast, Residual RL provides an efficient alternative to refine the base policy with minimal additional computation. This residual correction enables the agent to make targeted improvements, however the base policy was built.
Despite the promise of Residual RL, existing algorithms suffer from unconstrained exploration, often requiring extensive online interaction and dense reward shaping to achieve meaningful improvements [6, 7]. Furthermore, recent advancements in imitation learning leverage highly effective stochastic policies: Gaussian mixture model-based policies [8] and Diffusion policies [3] excel at modeling complex, multi-modal distributions. In such cases, residual RL algorithms that assume a deterministic base policy [1, 2] are not suitable.
We address these limitations by proposing two improvements to Residual RL that enhance its sample efficiency and make it more suitable for stochastic policies. First, we leverage uncertainty estimates from the base policy to guide the exploration of the residual policy. Our key insight is that regions where the base policy is confident require minimal exploration by the residual agent, allowing it to focus exploration on areas of high uncertainty. This targeted exploration significantly improves the sample efficiency of residual learning.
Second, existing off-policy Residual RL algorithms learn the $Q$ function only for the residual action $a_{r}$, i.e. $Q(s,a_{r})$, implicitly assuming that the base policy’s action can be inferred from the state $s$. This is insufficient when dealing with stochastic base policies, since they sample different actions given the same state. In the stochastic setting, the Residual RL agent cannot infer a single base action, making it difficult to learn a good residual action. Prior works have attempted to solve this by using learned bottleneck features of the base policy as a prior for residual learning  [9] or by augmenting the observed state with the base action for on-policy learning  [10]. We propose an asymmetric actor-critic approach, in which the critic learns the $Q$ function for the fully observed action executed in the environment, comprising both the base action and the residual action, while the actor learns partial residual actions only. This formulation ensures that information about the stochastic base actions is available to the $Q$ function while also making the critic invariant to the split between the residual and base action.
We evaluate our approach on a variety of manipulation tasks from Robosuite [8] and Franka Kitchen tasks from D4RL [11]. We test our approach on both Gaussian Mixture Model-based and Diffusion-based base policies. Finally, we compare our approach with a state-of-the-art finetuning method [12], a demo-augmented RL method [13], and other Residual RL methods [14, 1, 2]. Our proposed approach outperforms or is comparable to other baselines in all tasks. We also perform several ablation studies to test various aspects of our algorithm and also deploy the learned policies on a real robot to demonstrate sim-to-real transfer. Our proposed approach is visualized in Figure 1 with the main contributions summarized as follows:
  1. 1.
    We present a novel algorithm to accelerate Residual Reinforcement Learning using uncertainty estimates.
  2. 2.
    We modify off-policy Residual Reinforcement Learning to work with stochastic base policies using an asymmetric actor-critic approach.
  3. 3.
    We validate our method on robotic manipulation tasks from different simulators and against several baselines. We demonstrate zero-shot sim-to-real transfer of our learned policies.
II Background
In Residual RL, we assume that we have a suboptimal base policy $\pi_{b}$. The objective is to learn a lightweight residual policy $\pi_{r}$, on top of the base policy that gives a corrective action $a_{r}$ to produce a more accurate and robust policy. Thus, the final policy executed in the environment is :
$\pi=\pi_{b}(s)+\pi_{r}(s)$ (1)
Residual RL transforms the original Markov decision process (MDP) formulation $M=<S,A,R,T,\gamma>$ to the residual MDP (RMDP) formulation $M_{r}=<S,A_{r},R,T_{r},\gamma>$ [1], where $S$ is the set of states, $A_{r}$ is the set of residual actions, $R$ is the reward received for taking action $a_{r}\in A_{r}$ in state $s\in S$, $T_{r}$ is the probability of taking action $a_{r}$ in state $s$ and ending up in a new state $s^{\prime}$ and $\gamma$ is the discount factor. The residual transition function can be converted back to the original transition function as follows :
$T_{r}(s,a_{r},s^{\prime})=T(s,\pi_{b}(s)+a_{r},s^{\prime})\ .$ (2)
III Related Work
Residual RL for Stochastic Base Policies. Residual RL, first introduced for robotics in [2] and [1], learns a corrective residual policy over a base controller, which can be either hand-designed or derived from model-predictive control. Importantly, these methods assume a deterministic base controller, as the residual policy is not conditioned on the base action. However, current state-of-the-art imitation learning algorithms like Diffusion policy [3] and GMM-based policies [8] are non-deterministic, making the original Residual RL formulation insufficient due to the lack of information about the base policy. Some works introduce noise in the base action to enhance robustness and induce stochasticity [6, 7] . Other works inform residual learning about the base policy by conditioning it on the learned bottleneck features of the base policy [9], and incorporating the base action in the observed state to inform the residual policy for on-policy learning [10]. In contrast, our work modifies off-policy RL to learn the $Q$ function for the combined action taken in the environment (i.e. the sum of base and residual action), and the actor uses the same $Q$ function to select a residual action. Therefore, our Residual RL formulation can handle the stochasticity of the base policies by making the base action observable to the critic while also being invariant to the split between the base action and residual action. Policy Decorator [14] also uses the combined action as an input to their critic for Residual RL, but we do a thorough quantitative evaluation with ablations to show that it significantly improves performance.
Imitation Learning and Residual RL. Several works have explored the integration of base policies trained using Imitation Learning (IL) with Residual RL. Residual RL has been applied to insertion tasks [15], where demonstrations are incorporated as an auxiliary behavior cloning (BC) loss during RL training [16]. FISH [17] employs a non-parametric base policy alongside a residual policy that uses optimal transport matching against offline demonstrations as the reward. BeT [18] introduces a residual action corrector that refines continuous actions on top of a discretized imitation policy. IBRL [13] does not directly use Residual RL, but it bootstraps an RL policy from an IL policy by using IL actions as alternative proposals for both online exploration and critic updates. Closest to our method is Policy Decorator [14], which learns bounded residual actions using controlled exploration. Unlike Policy Decorator, which uniformly samples action from the base policy and residual policy, we use uncertainty estimates of the base policy to decide when to learn corrective residual actions.
Uncertainty Estimates in Imitation Learning. Uncertainty estimation plays a crucial role in improving the reliability and robustness of machine learning models, particularly in decision-making and RL. Various approaches have been proposed to quantify uncertainty, including distance-based techniques [19, 20], ensemble-based techniques [21, 22, 23], and learning another model to estimate uncertainty [24, 25, 26]. SGP [20] proposes a method that measures the distance of a given input to the training data distribution. This approach assumes that samples farther from the training distribution exhibit higher uncertainty, making it particularly useful for detecting out-of-distribution (OOD) inputs and improving model generalization. Diff-DAgger [27] uses the loss function of a Diffusion model to estimate uncertainty, where a higher loss indicates greater uncertainty. Our algorithm is agnostic to the uncertainty quantification method, and we test our approach with different uncertainty metrics.
IV Uncertainty aware Residual RL for stochastic policies
We describe how to incorporate Uncertainty Estimates in Residual RL in Sec. IV-A and our modified off-policy RL algorithm in Sec. IV-B.
IV-A Uncertainty Aware Residual RL
Prior works in Residual RL suffer from unrestrained exploration as they learn corrective residual actions uniformly over the entire state space. Our key insight is to improve exploration by focusing residual learning on regions in which the base policy is not confident. We propose using the uncertainty of the base policy to decide when to learn a residual action for the base policy. If the base policy is certain about its action for the current state, we directly use the action from the base policy $a_{b}$ in the environment; we only use a corrective residual action when the base policy is uncertain. Our proposed approach is agnostic to the uncertainty quantification method, and we demonstrate our method with two metrics, namely distance-to-data and ensemble variance. Distance-to-data has been used to calibrate the uncertainty of a model by measuring how out-of-distribution the current state is from an existing dataset [20]. For a dataset $D$ with each datapoint having $F$ features, we can estimate uncertainty using the minimum L2 distance of the current state $s$ to all points $d\in D$ :
$\text{uncertainty}_{d}(s)=\min_{d\in D}\sqrt{\sum_{i=1}^{F}(d_{i}-{\color[rgb]{0,0,0}s_{i}})^{2}}\ .$ (3)
Another popular approach to estimate uncertainty is by measuring the variance in predicted actions among an ensemble of policies. For an ensemble of $N$ base policies $\pi_{b}\in\pi_{B}$, the ensemble uncertainty can be defined as -
$\text{uncertainty}_{e}(s)=\frac{1}{N}\sum_{b=1}^{N}\left(\pi_{b}(a\mid s)-\frac{1}{N}\sum_{i=1}^{N}\pi_{i}(a\mid s)\right)^{2}\ .$ (4)
Furthermore, we use an uncertainty threshold $\tau$ to measure the confidence of the base policy. This can be formulated as
$a_{\text{taken}}=\begin{cases}a_{\text{b}}&\text{if uncertainty}<\tau\\ a_{\text{b}}+a_{\text{r}}&\text{otherwise}\ .\end{cases}$ (5)
As learning progresses, we decay this uncertainty threshold $\tau$ exponentially from a maximum uncertainty threshold value $U$ according to the following equation:
$\tau=U*e^{\frac{-\text{step}}{\text{decay\ rate}}}\ .$ (6)
The uncertainty threshold $\tau$ ultimately decays to 0 to let the residual policy take over. We perform ablations for different decaying strategies in Sec. V-G1.
IV-B Optimizing Residual RL for Stochastic Policies
The original Residual RL algorithms are formulated to learn only using the partial residual actions, operating under the assumption that the underlying base policy is deterministic and can be implicitly inferred. Therefore, it learns the $Q$ function for only the partial residual action which is different from the action taken in the environment:
$Q(s,a_{r})=\mathbb{E}_{s^{\prime}\sim P}\left[R(s,a_{r},s^{\prime})+\gamma V^{\pi}(s^{\prime})\right].$ (7)
Incorporating stochastic policies into the residual transition function $T_{r}$ can make it much harder to learn as they are noisier in their predictions and hence difficult to model. Thus, we suggest using Eq 2 to retreat back to the original MDP formulation. Consequently, the $Q$ function is learned for the combined action $a_{c}$, which is the actual action used during environment interaction :
$Q(s,a_{c})=\mathbb{E}_{s^{\prime}\sim P}\left[R(s,a_{b}+a_{r},s^{\prime})+\gamma V^{\pi}(s^{\prime})\right]$ (8)
Previously, ResiP [10] proposed augmenting the observed state with base action to provide the missing information for on-policy RL. We instead propose learning the critic for the combined action $a_{c}$ for off-policy RL to provide the necessary information about the base policy to the $Q$ function while also making it invariant to the split between residual action and the base action. Specifically, we modify Soft Actor-Critic [28] in the following ways (changes marked in green). Initially, we store both the base action $a_{b}$ and the combined action $a_{c}$ in the replay buffer. While computing the target values, we should add the base action $a_{b}$ to the residual action $a_{r}$ sampled from the actor as follows:
$\begin{split}y{\color[rgb]{0,0,0}(r,s^{\prime},d)}&=r+\gamma(1-d)\ *\\ &\quad(\min_{i=1,2}Q_{\phi^{\prime}_{i}}(s^{\prime},{\color[rgb]{0,.5,.5}a^{\prime}_{r}+{a^{\prime}_{b}}})-\alpha\log\pi_{r}({\color[rgb]{0,.5,.5}a^{\prime}_{r}}|s^{\prime})),\\ &\quad{\color[rgb]{0,.5,.5}a^{\prime}_{r}\sim\pi_{r}(\cdot|s^{\prime})}\ .\end{split}$ (9)
While updating the $Q$ function, we should use the combined action stored in the replay buffer:
$J_{Q}(\phi_{i})=\mathbb{E}\left[\left(Q_{\phi_{i}}(s,{\color[rgb]{0,.5,.5}a_{c}})-y{\color[rgb]{0,0,0}(r,s^{\prime},d)}\right)^{2}\right],\quad i=1,2\ .$ (10)
When updating the actor, we can again add the base action to get the $Q$ value:
$\begin{split}J_{\pi}(\theta)&=\mathbb{E}\left[Q_{\phi_{i}}(s,{\color[rgb]{0,.5,.5}a_{r}+a_{b}})-\alpha\log\pi_{r}({\color[rgb]{0,.5,.5}a_{r}}|s)\right],\\ &\quad i=1,2,{\color[rgb]{0,.5,.5}\quad a_{r}\sim\pi_{r}(\cdot|s)}\ .\end{split}$ (11)
The complete algorithm is described in Alg. 1 with the proposed changes in SAC marked in green.
Algorithm 1 Uncertainty aware Residual RL
1:Initialize parameters of residual policy $\pi_{r}$, Q-functions $Q_{\phi_{1}},Q_{\phi_{2}}$, and temperature $\alpha$
2:Initialize target Q-function parameters $\phi^{\prime}_{1}\leftarrow\phi_{1}$, $\phi^{\prime}_{2}\leftarrow\phi_{2}$
3:Initialize base policy $\pi_{b}$
4:for each environment step do
5:    Sample residual action from actor
6:           $a_{r}\sim\pi_{r}(s_{t})$
7:    Sample base action from base policy
8:           $a_{b}\sim\pi_{b}(s_{t})$
9:    Calculate uncertainty threshold $\tau$ for the base policy, Eq. 6
10:    Calculate the uncertainty in base policy, Eq. 3
11:    Select the action to be taken in the environment, Eq. 5
12:    Observe next state $s_{t+1}$ and reward $r_{t}$
13:    Store $(s_{t},a_{c},a_{b},r_{t},s_{t+1})$ in replay buffer $\mathcal{D}$
14:end for
15:for each gradient update step do
16:    Sample a minibatch $\{(s,a_{c},a_{b},r,s^{\prime})\}$ from $\mathcal{D}$
17:    Compute target value according to Eq 9
18:    Update Q-functions by minimizing according to Eq. 10
19:    Update policy $\pi_{\theta}$ using Eq. 11
20:    Update target networks $\phi^{\prime}_{i}$
21:end for
V Experiments
The setup for our experiments is described in Sec. V-A. We provide details on the baselines used in Sec. V-B and conduct an analysis of our results in Sec. V-C. We compare our approach with deterministic and stochastic base policy in Sec. V-D. We extend our approach to image-based environment in Sec V-E and perform real-world experiments in Sec V-F. Finally, we describe our ablation studies in Sec. V-G.
V-A Experiment Setup
We define the robotic manipulation tasks for our experiments in Sec. V-A1. All environments use sparse rewards and state-based observations with task visualizations in Fig. 2. We evaluate all experiments with 5 seeds over 20 runs and provide hyperparameters in Appendix VI-C.
V-A1 Environments
Robosuite: We evaluate the performance of the Panda robotic arm on three distinct tasks from the Robosuite simulator  [8]. We receive a sparse reward of 1 for successfully completing the task and 0 otherwise. We use the following tasks: 1) Lift task, where the robot arm must pick up a block placed at a random initial position on the table. 2) Can task, where the robot arm has to pick up a can from one table and place it in the top right corner of another table. 3) Square task, where the robot arm must pick up a square nut from the table and place it onto a square bolt.
Franka Kitchen: The Franka Kitchen environment from the D4RL benchmark [11] features a Franka robot that is required to interact with various objects to achieve a multitask goal configuration. It receives a reward of 1 for successfully completing each of the 4 sub-goals, and we report the normalized reward for each trajectory. The environment includes three datasets for the Franka Kitchen task from the D4RL benchmark: 1) Kitchen Complete: This dataset is limited in size and consists solely of positive demonstrations. We perform additional experiments with the other two datasets. 2) Kitchen Mixed: This dataset includes undirected demonstrations, with a portion of them successfully solving the task. 3) Kitchen Partial: This dataset also contains undirected demonstrations, but none of them fully solve the task. However, each demonstration successfully addresses certain components of the task.
V-A2 Base Policies
We consider two kinds of IL base policies to test the robustness of our algorithm :
Figure 2: We test our proposed approach on the Lift, Can, and Square tasks from Robosuite [8] and the Franka Kitchen Task from D4RL [11].
Figure 3: Results on Robosuite environments with a GMM base policy. Our method is able to outperform all other baselines in all tasks. The error bars indicate 95% confidence interval.
Figure 4: Results on Franka Kitchen and Robosuite environments with a Diffusion base policy. Our method is able to outperform all baselines for Kitchen Complete and Can task, and has comparable performance for Square Task. The error bars indicate 95% confidence interval.
Gaussian Mixture Model-based policy: We utilize GMM-based behavior cloning (BC) policies from [8], which has an RNN backbone. To introduce an additional challenge, we train policies for the Lift and Can tasks using noisier multi-human demonstrations. Since the Square task is inherently challenging, we use proficient-human demonstrations, as even high-quality demonstrations struggle to completely solve the task.
Diffusion-based policy: To ensure a consistent comparison, we use the same Diffusion policies from DPPO [12] for our experiments. For Robosuite tasks, they use noisier multi-human demonstrations, while for the Franka Kitchen environments, they directly use datasets from D4RL benchmark.
V-B Baselines
We compare our proposed approach against finetuning methods, demo augmented RL methods, and other Residual RL methods.
Finetuning methods: As a baseline for Diffusion-based policies, we use the recently proposed DPPO [12]. DPPO formulates the denoising process of the Diffusion policy as a separate MDP, effectively modeling the entire trajectory as a sequence of MDPs. The policy is then optimized using policy gradient across this entire chain of MDPs.
Demo augmented RL methods: We compare our method against two demo-augmented RL approaches. IBRL [13] maintains both an IL policy trained on demonstrations and an RL policy trained from scratch. During environment interaction and RL training, both policies propose actions, and the action with the highest $Q$ value is selected. A variant of this approach, IBRL-RPL, replaces the RL policy learned from scratch with a residual policy. We conduct experiments with both versions of IBRL.
Residual RL methods: We test our method against the most closely related approach and current state-of-the-art residual learning algorithm, Policy Decorator [14], which also aims to mitigate excessive exploration in Residual RL. It addresses this by sampling actions uniformly from both the residual and base policies while gradually decreasing the proportion of actions taken from the base policy. Additionally, it bounds the actions of the residual policy. For completeness, we also compare our approach with the standard Residual RL [1, 2] algorithm, incorporating our proposed modifications to the critic.
V-C Results and Analysis
We analyze the results for GMM-based policies in Sec. V-C1, and the results of Diffusion policies in Sec. V-C2. We compared the performance of our method with both distance-to-data and ensemble variance as the uncertainty metric for these results.
V-C1 GMM-based policies
We present the results of our experiments with GMM-based policies in Figure 3. We plot the success rate over the course of environment interactions. Our method with distance to data metric (red line) outperforms all the baselines in the three Robosuite environments. We note that there is an initial dip in the performance for our method where the residual policy is in the exploration phase, but it starts improving once the exploration phase ends. IBRL performs the best out of the other baselines with its performance comparable to our method for the Lift Task, though the initial dip in performance is more significant, and it is still unstable afterwards. We performed a hyperparameter sweep for the two additional parameters of Policy Decorator, namely the residual bound and the decay rate. Our method converges in the same number of timesteps while Policy Decorator’s performance is still improving. We suspect it is due to the more targeted exploration of our method using uncertainty estimates. The standard Residual Policy Learning method only goes over the base policy performance for the Lift Task.
V-C2 Diffusion policies
We present the results for Diffusion policies in Figure 4. We run the experiments for the Kitchen Complete environment of the D4RL benchmark. In Robosuite, we perform experiments in the Can and Square environments, excluding the Lift environment because of the near-optimal performance of Diffusion policy on that task. We performed a hyperparameter sweep of Policy Decorator as mentioned in Sec. V-C1. Similarly, our method is able to outperform Policy Decorator in all environments. Our approach is able to achieve higher success rates than all the baselines in the Kitchen Complete environment. We note that even though DPPO’s performance is stable in Kitchen Complete and Can environments, its improvement over the initial performance of the base policy is slow compared to our approach, despite an initial dip in the performance. Our method with ensemble variance as a metric has comparable performance to DPPO for the Square task. These results suggest that our approach is most promising in scenarios where the initial base policy performance is average, and is comparable in scenarios where the initial base policy performance is bad. Also, ensemble variance as a metric works better or is comparable to distance-to-data for the Can and Square environment but not in the Kitchen Complete environment. We hypothesize that this is because of the good quality of demonstrations in the Kitchen Complete environment while the noisier multi-human demonstrations of Can and Square tasks can result in a noisier distance-to-data metric, though our method is still able to converge for the same.
Figure 5: Learning with combined action and residual action for stochastic and deterministic base policies.
Figure 6: Results on Can-Image Task.
V-D Combined Action vs Residual Action
To emphasize the significance of utilizing combined actions for stochastic policies, we compared our modified SAC algorithm for Residual RL without uncertainty estimates to the original Residual RL formulation. For this comparison, we used the GMM policy as the stochastic base policy and a standard MLP policy as the deterministic base policy, applied to the Lift task in Robosuite. The results of using combined actions and residual actions during learning for both stochastic and deterministic base policies are shown in Figure 5. The findings reveal that relying solely on the residual action does not yield effective results for stochastic base policies, highlighting the necessity of combining actions in such cases. In contrast, for deterministic base policies, either residual actions or combined actions can be effectively used.
V-E Image Based Experiments
The results of our algorithm for the image-based Can Task is shown in Fig 6. Given that distance-to-data is a less reliable metric in high dimensional input spaces such as images, we instead used ensemble variance to quantify uncertainty. We explored two strategies: first, using ensemble variance to enhance the performance of a single policy, and second, leveraging it to improve an ensemble of policies. Our method was compared against DPPO, the strongest baseline for finetuning diffusion policies, and Residual Policy Learning (RPL) without uncertainty estimation. The results indicate that the ensemble-based approach starts with strong performance and avoids early dips, likely due to the robustness offered by ensembling. By contrast, the naive RPL strategy initially crashes to 0% performance after the initial online learning steps, though it recovers over time. DPPO, as observed previously, maintains steady but slow improvement. These findings show our approach also applies to image-based environments.
Figure 7: We present a novel approach for accelerating Residual RL. The learned policies were deployed in real world for the pick and place can task where the robot must pick the can from one side of the table and place it in the blue bin
Figure 8: Results for the policy deployment in real world. RL policies are robust in sim2real transfer even without any domain randomization.
V-F Real World Experiments
Additionally, we deployed our learned policy for the Can task in the real world in a zero-shot setup as shown in Fig 7. The task is setup similarly to the simulated version, where the robot has to pick up the can from one side of the table and place it in the blue bin on the other side of the table. We used Grounded SAM [29] and Foundation Pose [30] to get the real-time state of the can object. We evaluated four policies in our real-world setup: 1) GMM base, 2) GMM base + Residual, 3) Diffusion base, and 4) Diffusion base + Residual. We compare the performance in terms of task progression divided into three stages: contacting, grasping, and placing. We note the performance of each policy on 10 trials in Fig 8. The performance of policies learned with Residual RL is able to nearly retain their original performance in simulation without any domain randomization, whereas the base policies struggled. This supports the observation that policies trained using RL tend to be more robust than behavior cloning-based policies, as they benefit from richer interaction with the environment.
V-G Ablations
We ablated our proposed approach with different decaying strategies for the uncertainty threshold in Sec. V-G1, with different decay rates in Sec. V-G2, and for other kitchen environments in Sec. V-G3. All ablations are performed with the Diffusion base policy for the Can task.
Figure 9: Ablations for different threshold decay strategies. Decaying exponentially performs best.
V-G1 Threshold Decay Strategy
We tried different strategies for decaying the uncertainty threshold $\tau$: exponentially decaying the threshold to zero, exponentially decaying to a minimum threshold, and keeping the threshold constant. We plot success rate and the percentage of base policy actions used in Figure 9. We observe that exponential decay has the most stable performance out of the three. Exponentially decaying to a minimum threshold converges at a lower success rate compared to exponentially decaying the threshold to 0. The base actions used when decaying the threshold to a minimum value start increasing once that minimum threshold value is reached, signifying that the optimal policy stays within the distribution of the base policy. Keeping the threshold value constant restricts the residual policy from escaping the performance of the initial base policy.
Figure 10: Ablations for different decay rates.
Figure 11: Ablations for the other kitchen environments that contain undirected demonstrations.
V-G2 Threshold Decay Rate
We evaluated our algorithm with different decay rates in Figure 11. Lower rates resulted in aggressive exploration, causing performance dips that were hard to recover from, while higher rates slowed the convergence. We need to balance decay rates for effective exploration without hindering convergence. However, our approach is not too sensitive to the chosen decay rate, as the performance is similar for decay rates ranging between 300k and 500k.
V-G3 Kitchen Environments
We tested our method on the other two variations of the kitchen environments, Kitchen Partial and Kitchen Mixed. As seen in Figure 11, our method is not able to outperform the base policy. One key assumption we make in our algorithm is that when the base policy is certain, it will also be correct. This assumption does not hold true for these two environments: the base policies for these two environments are trained on the random play data, so the confidence and correctness of the policy are not strongly correlated, resulting in poorer performance.
VI Conclusion and Future Work
We propose two improvements to the Residual RL framework to accelerate the learning with stochastic base policies. First, we use uncertainty estimates of the base policy to constrain the exploration of residual learning. Second, we adapt Residual RL to handle stochastic base policies by proposing an asymmetric actor-critic approach in which the critic observes the combined action, while the actor predicts only the residual action. Our method outperforms various baselines on tasks from Robosuite and the D4RL Benchmark. While our proposed method demonstrates strong performance, it would also benefit from a more robust epistemic uncertainty metric. We believe that, with reliable uncertainty metrics, our approach could also be applied to larger models including robot foundation models.
Acknowledgments
This work was supported by NASA-Kennedy under grant GR5227163. SV was supported by the Office of Naval Research (ONR) under REPRISM MURI N000142412603 and ONR grant #N00014-22-1-2592, as well as by the National Science Foundation (NSF) via grant #1955361. Partial funding for SV was also provided by the Robotics and AI Institute. We would also like to thank Ondrej Biza for helpful discussions around the baselines and Ivy He for her help with real robot experiments.
APPENDIX
VI-A Environments
The observation space and the action space of each environment used are below -
  1. 1.
    Lift - 19 dim observation space with object state and robot end effector state. Action space is 7 DoF end effector pose.
  2. 2.
    Can - 23 dim observation space with object state and robot end effector state. Action space is 7 DoF end effector pose.
  3. 3.
    Square - 23 dim observation space with object state and robot end effector state. Action space is 7 DoF end effector pose.
  4. 4.
    Kitchen - 60 dim observation space with all object states and velocities, and robot joint states and angular velocity. The action space is 9 DoF joint angular velocity and gripper linear velocity.
VI-B Base Policies
VI-B1 GMM Base Policy
We use Robomimic [8] to train Gaussian Mixture Model based policies with a Recurrent Neural Network backbone.
VI-B2 Diffusion policy
We used the same base policies from DPPO [12] to keep the comparisons consistent. The Diffusion policies are trained with an action horizon of 1 and 20 denoising steps.
VI-C Hyperparameters
We used the same hyperparameters in each environment for both GMM-based and Diffusion-based. We keep the same parameters for actor and critic for Robosuite environments and use the advised hyperparameters from the DPPO paper for the kitchen environment. Hyperparameter details can be found in Table I. The uncertainty threshold value $U$ and decay rate values for our proposed approach with distance-to-data metric can be found in Table II, and with ensemble variance can be found in Table III.
Environment Actor & Critic Dimensions Actor lr Critic lr
Robosuite (256,256) 1e-4 1e-4
Kitchen (256,256,256) 1e-5 1e-3
TABLE I: Hyperparameters used for each environment
Environment Base Policy U Decay Rate
Lift GMM 1e-6 200k
Can GMM 2e-5 75k
Square GMM 5e-5 150k
Kitchen Complete Diffusion 2.5e-3 200k
Can Diffusion 4.5e-5 400k
Square Diffusion 4.5e-5 1M
TABLE II: Uncertainty threshold $U$ and Decay Rate values for distance-to-data method
Environment Base Policy U Decay Rate
Kitchen Complete Diffusion 0.5 200k
Can Diffusion 0.2 500k
Square Diffusion 0.4 750k
TABLE III: Uncertainty threshold $U$ and Decay Rate values for ensemble variance
VI-D Tuning Policy Decorator
Policy Decorator  [14] has two hyperparameters namely, $\alpha$ the residual bound which scales the residual action to limit exploration and $H$ to schedule the exploration progressively. According to their paper, the $\alpha$ value is set close to the action scale of demonstration data while it is advised to keep $H$ large as a safe choice. We present the hyperparameter values we used in our sweep in Table IV. The authors perform an ablation with DPPO for the square task in their appendix, and we received the hyperparameters from the authors for the same. After looking at their implementation, we observed that they train their RL policies with an expanded action space with an action horizon while we implement the RL policies in the original form with a single action.
Environment Base Policy $\boldsymbol{\alpha}$ H
Lift GMM 0.1, 0.2, 0.05 400k, 600k
Can GMM 0.05,0.1,0.2,0.5 400k, 600k, 800k
Square GMM 0.05, 0.1, 0.5 750k, 1M
Kitchen Complete Diffusion 0.1, 0.2, 0.3 400k, 600k
Can Diffusion 0.2, 0.5 400k
Square Diffusion 0.05 500k
TABLE IV: Residual bound $\alpha$ and Progressive Exploration schedule $H$ for Policy Decorator